OpenStudy (marcelie):

help please find a power series representation for the function and determine the interval of convergence

OpenStudy (marcelie):



OpenStudy (marcelie):

@zepdrix

zepdrix (zepdrix):

you -_-

zepdrix (zepdrix):

Ummm

OpenStudy (marcelie):

lool

zepdrix (zepdrix):

Let's factor a 3 out of each term in the denominator, and the 2 out of the top,\[\large\rm =\left(\frac23\right)\frac{1}{1-\frac x3}\]

zepdrix (zepdrix):

Does it look like anything familiar in this form? Hmmm

OpenStudy (marcelie):

aaaah wait what formula we using lol gosh its been weeks i forgot about this :O

zepdrix (zepdrix):

I'll give a hint :O\[\large\rm \frac{1}{1-r}\]

OpenStudy (marcelie):

ohh that formula